OpenStudy (anonymous):

solve for x 4x^2=-4x+3

OpenStudy (anonymous):

Firstly bring all terms to one side.. You can do by adding 4x both the sides and then subtracting 3 from both the sides.. You will get: \(4x^2 + 4x - 3 = 0\)

OpenStudy (anonymous):

\[4x^2 + 4x - 3 = 0 \implies 4x^2 + 6x-4x-3 = 0 \implies 2x(2x+3) - 1(2x+3) = 0 \\ \quad \quad \quad \quad \quad \quad \quad \implies \color{green}{(2x-1)(2x+3) = 0}\]

OpenStudy (anonymous):

Now, either first bracket will be 0 or second bracket or both.. So: Set the two brackets separately to \(0\), as find \(x\), you will get two values of \(x\)..